1.Reconsider the example above, where you are contemplating the purchase of the coupon bond with a face value of $1,000, which matures in 14 years, and pays 3.15% (annual) coupons. Now, If you require a return of 3.75% on this instrument, how much would you offer to pay for it today? [Present the answer rounded to two decimal places. For instance 1045.16]

2. You are considering the purchase of a Pure Discount Bond with a Face Value of $10,000, which matures in three years. If you desire a return of 3.35%, how much would you bid for the bond today? (Round your answer to two decimal places, e.g. 9274.36) (answer was not 9058.76)

Thank you!
Answer #1

1.
Using financial calculator
FV=-1000
N=14
PMT=-3.15%*1000=-31.5
I/Y=3.75%
CPT PV=935.56

2.
Using financial calculator
FV=-10000
N=3*2
PMT=0
I/Y=3.35%/2
CPT PV=9051.38